A triploid plant has 18 chromosomes (i.e., 6 chromosomes per set).
a. If we assume a gamete has an equal probability of receiving one or two copies of each of the six types of chromosome, what are the odds of this plant producing a monoploid or a diploid gamete?

b. What are the odds of producing an aneuploid gamete?
Answer #1

Given,

number of chromosomes per set (n) = 6

a)the odds of producing monoploid or diploid gamete = (1/2)n-1 = (1/2)6-1 = (1/2)5 = 1/32

Hence 1 out of 32 has the chance of producing monoploid or diploid gamete.

b) The odds of producing an aneuploid gamete is 31 out of 32.